OpenStudy (anonymous):

Please help. Write a rule for the nth term of the sequence. Use your rule to find a100.–8, 9, 26, 43, 60, . . .

OpenStudy (anonymous):

Looks like in A.P with first term -8 and common difference 17

OpenStudy (anonymous):

tn = t1 + (n-1)d where t1 is the first tem and d is the common difference

OpenStudy (anonymous):

Does it help?

OpenStudy (anonymous):

Kind of, but what does it mean when it says to write a rule? What's the rule?

OpenStudy (anonymous):

tn=-8+(n-1)17=-8+17n-17=-25+17n

OpenStudy (anonymous):

Thats the rule

OpenStudy (anonymous):

Now, u need to find t100 = -25+17*100

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

Yeah, thanks.